66. A survey of undergraduate students in the School of Business at Northern University revealed the following regarding the gender and majors of the students:
Major
Gender Accounting Management Finance Total
Male 100 150 50 300
Female 100 50 50 200
Total 200 200 100 500
a. What is the probability of selecting a female student?
b. What is the probability of selecting a finance or accounting major?
c. What is the probability of selecting a female or an accounting major? Which rule of addition
did you apply?
d. Are gender and major independent? Why?
e. What is the probability of selecting an accounting major, given that the person selected
is a male?
f. Suppose two students are selected randomly to attend a lunch with the president of the
university. What is the probability that both of those selected are accounting majors?

Answer by stanbon(48568)

A survey of undergraduate students in the School of Business at Northern University revealed the following regarding the gender and majors of the students:
Major
Gender Accounting Management Finance Total
Male.... 100........ 150...... 50.... 300
Female.. 100......... 50...... 50.... 200
Total... 200........ 200..... 100.... 500
-----------------------------------------------
a. What is the prob of selecting a female student?: 200/500 = 2/5
-------------------------------------------
b. What is the prob of selecting a finance or accounting major?: = 300/500
---------------------------------------------
c. What is the prob of selecting a female or an accounting major? Which rule of addition did you apply?: (200+200-100)/500= 3/5; P(AorB)=P(A)+P(B)-P(AandB)
-----------------------------------------------
d. Are gender and major independent? Why?: No because there is an intersection.
or P(male|acct) is not equal to P(male)*P(acct)
P(male|acct) = 1/2 ; P(male)*P(acct) = (3/5)(2/5) = 6/25
------------------------------------------------------------
e. What is the prob of selecting an accounting major, given that the person selected is a male? P(acct|male) = 100/200 = 1/2
------------------------------------------------------
f. Suppose two students are selected randomly to attend a lunch with the president of the university. What is the prob that both of those selected are accounting majors? : 200C2/500C2 = 0.15952...
-------------------------------------------------------
Cheers,
Stan H.